a. The market for vanilla ice cream is given by the following information:

๐‘„

๐‘‘ = 800 โˆ’ 30๐‘ƒ

๐‘ฃ + 10๐‘ƒ

๐‘

๐‘„

๐‘  = 250 + 30๐‘ƒ

๐‘ฃ โˆ’ 10๐‘ƒ๐‘š

Where ๐‘„

๐‘‘

is the quantity demanded, ๐‘„

๐‘ 

is quantity supplied, ๐‘ƒ

๐‘ฃ

is the price of vanilla iceย 

cream, ๐‘ƒ

๐‘

is the price of chocolate ice cream and ๐‘ƒ๐‘š is the price of milk.


1
Expert's answer
2021-03-22T11:57:44-0400

At equilibrium, supply is equal to demand. So;

"Qd=Qs"

"(800-30Pv+10Pc)=(250+30Pv-10Pm)"

We know thatย "Pc=10"ย andย "Pm=5"

hence;ย "800-30Pv+10(10)=250+30Pv-10(5)"

"30Pv+30Pv=800+100+50-250"

"Pv=12"ย , which the price of Vanilla ice cream at equilibrium.

Therefore, quantity supplied at equilibrium isย "250+30(12)-10(5)"

"=560"


The quantity demanded at equilibrium will beย "800-30(12)+10(10)"

"=540"